Subtract.4 1/3-1 1/4
Answer by Edwin McCravy(20086)


4 1/3-1 1/4 

Write them vertically like this

 4 1/3
-1 1/4 


Get the LCD of 3 and 4. Since they have no common
factors other than 1, the LCD is their product, 12.

To start, we need to change the 1/3 to something 
over 12.

Notice that the denominator of 1/3, which is 3,
needs to be multiplied by 4 to become 12. So 
multiply both the denominator and numerator by 4.
That gives you a numerator of 4 and a denominator 
of 12.  So put an equal sign after 4 1/3 and put 
4 4/12 after it, like this


  4 1/3 =  4 4/12  
 -1 1/4   
          
Now we need to change 1/4 to something over 12.

Notice that the denominator of 1/4, which is 4,
needs to be multiplied by 3 to become 12. So 
multiply both the denominator and numerator by 3.
That gives you a numerator of 3 and a denominator 
of 12.  So put an equal sign after -1 1/4 and put 
1 3/12 after it, like this 

  4 1/3 =   4 4/12  
 -1 1/4 =  -1 3/12

Now we draw a line underneath like this

  4 1/3 =   4 4/12  
 -1 1/4 =  -1 3/12
          ---------

Now we subtract the whole parts: 4 - 1 = 3
and we subtract only the numerators of the
fractions, 4-1, getting a numerator of 3,
keeping the denominator of 12, like this:

  4 1/3 =   4 4/12  
 -1 1/4 =  -1 3/12
          ---------
            3 1/12

When the numerator of the top fraction is
smaller than numerator of the bottom
fraction, you have to do something else,
but that wasn't required here.

Edwin